# Model Predictive Inferential Control of Neural State-Space Models for Autonomous Vehicle Motion Planning

### v4

@@ V-D MPIC via Unscented Kalman-IPF and Kalman-IPS
$[\bm{m}^{\bm{y}},\bm{P}^{\bm{y}},\bm{P}^{\bm{x}\bm{y}}]$ $=\mathcal{UT}(\gamma,\bm{m}^{\bm{x}},\bm{P}^{\bm{x}},\bm{Q}),$
$\bm{y}$ $\sim\mathcal{N}(\bm{m}^{\bm{y}},\bm{P}^{\bm{y}}).$
@@ VI Discussion
Implementing the MPIC-X algorithm requires moderate parameter tuning to maximize its performance, mostly to ensure wide enough sampling ranges within the state and control space. The tuning effort hence is largely directed towards adjusting some covariance matrices used in the algorithm. First, $\bar{\bm{Q}}$ and $\bar{\bm{R}}$ may need to be inflated, as aforementioned. Second, for the horizon $[k,k+H]$, ${\bm{\Sigma}}_{k}^{f}=\mathrm{diag}\left(\bm{0},\bm{\Sigma}^{f,\bm{u}}_{k},\bm{\Sigma}_{k}^{f,\Delta\bm{u}}\right)$, the initial filtering covariance at time $k$, should be chosen such that the initial particles for $\bm{u}_{k}$ and $\Delta\bm{u}_{k}$ are explorative enough. Third, to draw highly probable particles for $\bm{\xi}_{t}$, we can sample from $\mathcal{N}\left(0,\bm{\Sigma}_{\mathrm{IIS}}\right)$, where $\bm{\Sigma}_{\mathrm{IIS}}=\mathrm{diag}(\sigma_{1}\bm{I},\sigma_{2}\bm{I},\sigma_{3}\bm{I})$ and $\sigma_{i}\leq 1$ for $i=1,2,3$, so that the particles are highly probable with respect to $\mathcal{N}\left(\bm{0},\bm{I}\right)$. Here, $\sigma_{i}$ for $i=1,2,3$ correspond to the implicit importance sampling for $\bm{x}_{k}$, $\bm{u}_{k}$, and $\Delta\bm{u}_{k}$, respectively, with $\sigma_{2},\sigma_{3}\gg\sigma_{1}$ for the sake of explorative control input. Finally, $\mathcal{UT}$ involves some tunable parameters, among which $\alpha$ adjusts the spread of the sigma points. The literature often suggests $\alpha=10^{-3}$ [8], but a much larger $\alpha$ is more effective here to enhance exploration due to the need of dealing with the highly nonlinear NSS model.
@@ VII-A The Simulation Setting
  • Net-$1$: a single hidden layer with $512$ neurons;
  • Net-$2$: two hidden layers with $128$ neurons in each layer;
  • Net-$3$: four hidden layers with $64$, $128$, $128$, and $64$ neurons in each layer.
@@ VII-B The Overtaking Scenario
Network Horizon ($H$) Method Total Cost Average Computation Time (s) Relative Cost Change w.r.t IPOPT & fmincon (%) Relative Computation Time Change w.r.t IPOPT & fmincon (%)
Net-$1$ $\qquad\;\,{\scriptstyle(512)}$ $10$ CasADi w. IPOPT $16,808$ $0.159$
@@
MPIC-X w. $10$ particles $18,574$ $0.025$ $10.51$  &  $3.64$ $-84.28$  &  $-99.69$
MPIC-X w. $50$ particles $18,158$ $0.069$ $8.04$    &  $1.32$ $-56.60$  &  $-99.15$
MPIC-X w. $80$ particles $18,077$ $0.096$ $7.56$    &  $0.86$ $-39.62$  &  $-98.82$
$20$ CasADi w. IPOPT $63,197$ $12.58$
@@
MPIC-X w. $10$ particles $18,235$ $0.052$ $-71.15$  &  $-3.33$ $-99.59$  &  $-99.78$
MPIC-X w. $50$ particles $17,701$ $0.147$ $-71.99$  &  $-6.16$ $-98.83$  &  $-99.37$
MPIC-X w. $80$ particles $17,602$ $0.202$ $-72.15$  &  $-6.68$ $-98.40$  &  $-99.15$
$40$ CasADi w. IPOPT
@@
MPIC-X w. $10$ particles $18,596$ $0.105$ —  &  $-38.22$ —  &  $-99.80$
MPIC-X w. $50$ particles $17,944$ $0.303$ —  &  $-40.39$ —  &  $-99.43$
MPIC-X w. $80$ particles $17,890$ $0.415$ —  &  $-40.57$ —  &  $-99.22$
$60$ CasADi w. IPOPT
@@
Net-$2$ $\quad\;\,{\scriptstyle(128-128)}$ $10$ CasADi w. IPOPT $16,650$ $0.211$
@@
[…]
MPIC-X w. $50$ particles $18,923$ $0.065$ $13.65$ $-0.48$ $-69.10$  &  $-99.12$
MPIC-X w. $80$ particles $18,778$ $0.091$ $12.77$ $-0.29$ $-56.82$  &  $-98.72$
$20$ CasADi w. IPOPT $19,499$ $0.337$
@@
MPIC-X w. $10$ particles $18,064$ $0.055$ $-7.36$ $4.97$ $-83.40$  &  $-99.75$
MPIC-X w. $50$ particles $17,633$ $0.141$ $-9.57$ $2.46$ $-58.04$  &  $-99.37$
MPIC-X w. $80$ particles $17,473$ $0.196$ $-10.4$ $1.53$ $-41.74$  &  $-99.13$
$40$ CasADi w. IPOPT
@@
Net-$3$ ${\scriptstyle(64-128-128-64)}$ $10$ CasADi w. IPOPT $16,396$ $0.380$
@@
MPIC-X w. $10$ particles $18,551$ $0.031$ $13.14$  &  $8.57$ $-91.84$  &  $-99.92$
MPIC-X w. $50$ particles $18,047$ $0.072$ $10.07$  &  $5.62$ $-81.05$  &  $-99.82$
[…]
$20$ CasADi w. IPOPT $26,088$ $2.163$
@@
MPIC-X w. $10$ particles $17,434$ $0.058$ $-33.17$  &  $12.13$ $-99.59$  &  $-99.78$
[…]
$40$ CasADi w. IPOPT
@@
Next, we compare the MPIC-X algorithm against gradient-based MPC in terms of computation time and planning performance. To solve gradient-based MPC, we use two solvers: MATLAB’s fmincon with the default interior point method, and CasADi [74] with interior point optimizer (IPOPT) [75].
For a fair comparison, the closed form of each gradient is pre-determined offline and then called during the online optimization process, and the optimization is also warm-started. The stopping tolerances for fmincon and CasADi-IPOPT are set close to each other to ensure impartiality in performance evaluation. To achieve a comprehensive assessment, we perform the simulation for different planning horizons by letting $H=10$, $20$, $40$, and $60$, and for the MPIC-X algorithm, we use different particle numbers, $N=10$, $50$, and $80$. Each setting comes with ten simulation runs to find out the average computation time.
Table VII-B summarizes the simulation results and comparisons. The overarching observations are as follows.
Figure 5: Cost performance comparison between the MPIC-X algorithm with ten particles and the gradient-based MPC for Net-1 at different horizon lengths: (a) $H=10$; (b) $H=20$; (c) $H=40$; (d) $H=60$. The red-shaded region represents ±$\pm 2\sigma$ bounds for ten Monte Carlo runs of the MPIC-X algorithm.
@@
  • The MPIC-X algorithm demonstrates strong capabilities in controlling NSS models. While IPOPT is much faster than fmincon as shown in Table VII-B, the MPIC-X algorithm outpaces IPOPT by a substantial degree. Using only $N=10$ particles, it reduces computation time by at least $80\%$ compared to IPOPT. Although the computation unsurprisingly increases as $N$ grows for the MPIC-X algorithm, the increase rates are less than linear, and the algorithm always outperforms IPOPT and fmincon, by at least $40\%$. It is important to emphasize the success of the MPIC-X algorithm in addressing motion planning even when the planning horizon $H=40$ and $60$, where IPOPT and fmincon fail to converge. Another appealing feature of the MPIC-X algorithm is its computational insensitivity to the neural network architecture. Due to its sampling-based nature, the algorithm achieves nearly indistinguishable increase in computation time for Net-1/2/3, making it highly advantageous for controlling NSS models with varying levels of neural network complexity. Further, the MPIC-X algorithm generates slightly higher, yet still close, costs than IPOPT and fmincon in the cases when the latter manage to converge. The simulations constantly show that the MPIC-X algorithm with just $N=10$ particles can deliver good enough cost performance and outstanding computational efficiency.
@@ VIII Motion Planning for a Real-World Vehicle
Next, the MPIC-X algorithm is applied to the trained NSS model to perform motion planning for overtaking. Fig. 9 illustrates that the red EV successfully overtakes the blue and green OVs with two lane changes. Figs. 9(a)-9(b) present the actuation profiles, while Fig. 9(c) depicts the distances between the EV and the OVs. Initially driving behind the green OV, the EV begins with mild acceleration and executes a lane change. Once a safe distance is achieved relative to both OVs, the EV steers back into the original lane, overtaking the green OV. The EV maintains the compliance with the actuation and safety constraints throughout the maneuvers. Table II offers a quantitative performance evaluation. Overall, the MPIC-X algorithm demonstrates high computational efficiency and scalability with respect to both the prediction horizon length and the number of particles. The algorithm’s cost performance is also excellent, with just ten particles sufficient to achieve low enough costs at fast computation. These results are consistent with what Table VII-B shows when considering the complexity of the neural network architecture, showing the potential effectiveness of the MPIC-X algorithm for real-world problems.
@@
Horizon ($H$) Method Total Cost Average Computation Time (s)

### v5

@@ V-D MPIC via Unscented Kalman-IPF and Kalman-IPS
$\left[\bm{m}^{\bm{y}},\bm{P}^{\bm{y}},\bm{P}^{\bm{x}\bm{y}}\right]$ $=\mathcal{UT}\left(\gamma,\bm{m}^{\bm{x}},\bm{P}^{\bm{x}},\bm{Q}\right),$
$\bm{y}$ $\sim\mathcal{N}\left(\bm{m}^{\bm{y}},\bm{P}^{\bm{y}}\right).$
@@ VI Discussion
Implementing the MPIC-X algorithm requires moderate parameter tuning to maximize its performance, mostly to ensure wide enough sampling ranges within the state and control space. The tuning effort hence is largely directed towards adjusting some covariance matrices used in the algorithm. First, $\bar{\bm{Q}}$ and $\bar{\bm{R}}$ may need to be inflated, as aforementioned. Second, for the horizon $[k,k+H]$, ${\bm{\Sigma}}_{k}^{f}=\mathrm{diag}\left(\bm{0},\bm{\Sigma}^{f,\bm{u}}_{k},\bm{\Sigma}_{k}^{f,\Delta\bm{u}}\right)$, the initial filtering covariance at time $k$, should be chosen such that the initial particles for $\bm{u}_{k}$ and $\Delta\bm{u}_{k}$ are explorative enough. Third, to draw highly probable particles for $\bm{\xi}_{t}$, we can sample from $\mathcal{N}\left(0,\bm{\Sigma}_{\mathrm{IIS}}\right)$, where $\bm{\Sigma}_{\mathrm{IIS}}=\mathrm{diag}\left(\sigma_{1}\bm{I},\sigma_{2}\bm{I},\sigma_{3}\bm{I}\right)$ and $\sigma_{i}\leq 1$ for $i=1,2,3$, so that the particles are highly probable with respect to $\mathcal{N}\left(\bm{0},\bm{I}\right)$. Here, $\sigma_{i}$ for $i=1,2,3$ correspond to the implicit importance sampling for $\bm{x}_{k}$, $\bm{u}_{k}$, and $\Delta\bm{u}_{k}$, respectively, with $\sigma_{2},\sigma_{3}\gg\sigma_{1}$ for the sake of explorative control input. Finally, $\mathcal{UT}$ involves some tunable parameters, among which $\alpha$ adjusts the spread of the sigma points. The literature often suggests $\alpha=10^{-3}$ [8], but a much larger $\alpha$ is more effective here to enhance exploration due to the need of dealing with the highly nonlinear NSS model.
@@ VII-A The Simulation Setting
  • Net-11: a single hidden layer with $512$ neurons;
  • Net-22: two hidden layers with $128$ neurons in each layer;
  • Net-33: four hidden layers with $64$, $128$, $128$, and $64$ neurons in each layer.
@@ VII-B The Overtaking Scenario
Network
Horizon
($H$)
Method Total Cost
Average
Computation
Time (s)
Relative Cost
Change w.r.t
IPOPT & fmincon (%)
Relative Computation
Time Change w.r.t
IPOPT & fmincon (%)
Net-11
$\qquad\;\,{\scriptstyle(512)}$
$10$
CasADi w. IPOPT $16,808$ $0.159$
@@
MPIC-X w. $10$ particles $18,574$ $0.025$ $10.51$  &  $3.64$ 84.28-84.28  &  99.69-99.69
MPIC-X w. $50$ particles $18,158$ $0.069$ $8.04$    &  $1.32$ 56.60-56.60  &  99.15-99.15
MPIC-X w. $80$ particles $18,077$ $0.096$ $7.56$    &  $0.86$ 39.62-39.62  &  98.82-98.82
2020
CasADi w. IPOPT $63,197$ $12.58$
@@
MPIC-X w. $10$ particles $18,235$ $0.052$ 71.15-71.15  &  3.33-3.33 99.59-99.59  &  99.78-99.78
MPIC-X w. $50$ particles $17,701$ $0.147$ 71.99-71.99  &  6.16-6.16 98.83-98.83  &  99.37-99.37
MPIC-X w. $80$ particles $17,602$ $0.202$ 72.15-72.15  &  6.68-6.68 98.40-98.40  &  99.15-99.15
$40$
CasADi w. IPOPT
@@
MPIC-X w. $10$ particles $18,596$ $0.105$ —  &  38.22-38.22 —  &  99.80-99.80
MPIC-X w. $50$ particles $17,944$ $0.303$ —  &  40.39-40.39 —  &  99.43-99.43
MPIC-X w. $80$ particles $17,890$ $0.415$ —  &  40.57-40.57 —  &  99.22-99.22
$60$
CasADi w. IPOPT
@@
Net-22
$\quad\;\,{\scriptstyle(128-128)}$
$10$
CasADi w. IPOPT $16,650$ $0.211$
@@
MPIC-X w. $10$ particles $18,992$ $0.029$ $17.31$ & $\phantom{-}3.71$ 86.26-86.26  &  99.59-99.59
MPIC-X w. $50$ particles $18,923$ $0.065$ $13.65$ & 0.48-0.48 69.10-69.10  &  99.12-99.12
MPIC-X w. $80$ particles $18,778$ $0.091$ $12.77$ & 0.29-0.29 56.82-56.82  &  98.72-98.72
2020
CasADi w. IPOPT $19,499$ $0.337$
@@
MPIC-X w. $10$ particles $18,064$ $0.055$ 7.36-7.36 & $4.97$ 83.40-83.40  &  99.75-99.75
MPIC-X w. $50$ particles $17,633$ $0.141$ 9.57-9.57 & $2.46$ 58.04-58.04  &  99.37-99.37
MPIC-X w. $80$ particles $17,473$ $0.196$ 10.4-10.4 & $1.53$ 41.74-41.74  &  99.13-99.13
$40$
CasADi w. IPOPT
@@
Net-33
$\scriptstyle(64-128-128-64)$}
$10$
CasADi w. IPOPT $16,396$ $0.380$
@@
MPIC-X w. $10$ particles $18,551$ $0.031$ $13.14$  &  $8.57$ 91.84-91.84  &  99.92-99.92
MPIC-X w. $50$ particles $18,047$ $0.072$ $10.07$  &  $5.62$ 81.05-81.05  &  99.82-99.82
MPIC-X w. $80$ particles $18,017$ $0.101$ $9.89\phantom{0}$  &  $5.45$ $73.4\phantom{0}  &$  99.75-99.75
2020
CasADi w. IPOPT $26,088$ $2.163$
@@
MPIC-X w. $10$ particles $17,434$ $0.058$ 33.17-33.17  &  $12.13$ 99.59-99.59  &  99.78-99.78
MPIC-X w. $50$ particles $16,970$ $0.152$ 34.95-34.95  &  $9.14\phantom{0}$ 98.83-98.83  &  99.37-99.37
MPIC-X w. $80$ particles $16,853$ $0.202$ 35.40-35.40  &  $8.39\phantom{0}$ 98.40-98.40  &  99.15-99.15
$40$
CasADi w. IPOPT
@@
Next, we compare the MPIC-X algorithm against gradient-based MPC in terms of computation time and planning performance. To solve gradient-based MPC, we use two solvers: MATLAB’s fmincon with the default interior point method, and CasADi [73] with IPOPT [74].
For a fair comparison, the closed form of each gradient is pre-determined offline and then called during the online optimization process, and the optimization is also warm-started. The stopping tolerances for fmincon and CasADi-IPOPT are set close to each other to ensure impartiality in performance evaluation. To achieve a comprehensive assessment, we perform the simulation for different planning horizons by letting $H=10$, 2020, $40$, and $60$, and for the MPIC-X algorithm, we use different particle numbers, $N=10$, $50$, and $80$. Each setting comes with ten simulation runs to find out the average computation time.
Table I summarizes the simulation results and comparisons. The overarching observations are as follows.
Figure 5: Cost performance comparison between the MPIC-X algorithm with ten particles and the gradient-based MPC for Net-1 at different horizon lengths: (a) $H=10$; (b) $H=20$; (c) $H=40$; (d) $H=60$. The red-shaded region represents $\pm 2\sigma$ bounds for ten Monte Carlo runs of the MPIC-X algorithm.
@@
  • The MPIC-X algorithm demonstrates strong capabilities in controlling NSS models. While IPOPT is much faster than fmincon as shown in Table I, the MPIC-X algorithm outpaces IPOPT by a substantial degree. Using only $N=10$ particles, it reduces computation time by at least $80\%$ compared to IPOPT. Although the computation unsurprisingly increases as $N$ grows for the MPIC-X algorithm, the increase rates are less than linear, and the algorithm always outperforms IPOPT and fmincon, by at least $40\%$. It is important to emphasize the success of the MPIC-X algorithm in addressing motion planning even when the planning horizon $H=40$ and $60$, where IPOPT and fmincon fail to converge. Another appealing feature of the MPIC-X algorithm is its computational insensitivity to the neural network architecture. Due to its sampling-based nature, the algorithm achieves nearly indistinguishable increase in computation time for Net-1/2/3, making it highly advantageous for controlling NSS models with varying levels of neural network complexity. Further, the MPIC-X algorithm generates slightly higher, yet still close, costs than IPOPT and fmincon in the cases when the latter manage to converge. The simulations constantly show that the MPIC-X algorithm with just $N=10$ particles can deliver good enough cost performance and outstanding computational efficiency.
@@ VIII Motion Planning for a Real-World Vehicle
Next, the MPIC-X algorithm is applied to the trained NSS model to perform motion planning for overtaking. Fig. 9 illustrates that the red EV successfully overtakes the blue and green OVs with two lane changes. Figs. 10(a)-10(b) present the actuation profiles, while Fig. 10(c) depicts the distances between the EV and the OVs. Initially driving behind the green OV, the EV begins with mild acceleration and executes a lane change. Once a safe distance is achieved relative to both OVs, the EV steers back into the original lane, overtaking the green OV. The EV maintains the compliance with the actuation and safety constraints throughout the maneuvers. Table II offers a quantitative performance evaluation. Overall, the MPIC-X algorithm demonstrates high computational efficiency and scalability with respect to both the prediction horizon length and the number of particles. The algorithm’s cost performance is also excellent, with just ten particles sufficient to achieve low enough costs at fast computation. These results are consistent with what Table I shows when considering the complexity of the neural network architecture, showing the potential effectiveness of the MPIC-X algorithm for real-world problems.
@@
Horizon ($H$)
Method Total Cost
Average
Computation
Time (s)